OpenStudy (anonymous):

How would you find the bounds for this polar integral Use a double integral to find the area of the region enclosed by the curve r=4+3cos(theta)

OpenStudy (anonymous):

\[r=4+3\cos \theta\]

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi}\int\limits_{0}^{4+3\cos \theta}\]

OpenStudy (anonymous):

how did you get 0 to pie for the interval of theta

OpenStudy (anonymous):

You didn't give me anymore info, I don't know if there is more. Changing r to rectangular I it is a half circle, 0 to pi.

OpenStudy (anonymous):

that is the question given in the book, there is nomore information

OpenStudy (anonymous):

you don't need more. If you graph it you'll know. It's a cardiod, so the limits are 0 to 2pi, if I remember correctly

OpenStudy (anonymous):

and the thing you would be integrating is r dr dtheta don't forget the r!

OpenStudy (anonymous):

would the integration for r still be from 0 to 4+3costheta

OpenStudy (anonymous):

That one is set, I just wasn't sure of the angle.

OpenStudy (anonymous):

yeah, i agree

OpenStudy (anonymous):

chaguanas, so i like posted my question 3 times and that guy you were talking about still hasn't shown up yet

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

I think she lives in Australia or something and comes on in the middle of the night, with luck she would reply to it.